Use Lagrange multipliers to find the maximum and minimum values of the function ( if they exist) subject to the specified constraint.

f(x,y) = 4x^2 + 9y^2 on the hyperbola xy = 6


1
Expert's answer
2022-02-01T15:41:03-0500

"f_x=8x,f_y=18y"

"g_x=y,g_y=x"


"f_x=\lambda g_x \implies 8x=\lambda y"

"f_y=\lambda g_y \implies 18y=\lambda x"


"\frac{4x}{9y}=\frac{y}{x}\implies 9y^2=4x^2"


"y=\pm 2x/3"


then:

"2x^2/3=6\implies x=\pm 3"


solution:

"(3,2),(3,-2),(-3,2),(-3,-2)"

"f(3,2)=f(3,-2)=f(-3,2)=f(-3,-2)=72"


"f(x,y) = 4(6\/y)^2 + 9y^2"

if we take, for example, y =1we get:

"f(x,1)=153>72"

so,

f(3,2)=f(3,-2)=f(-3,2)=f(-3,-2)=72 are minimums